OpenStudy (anonymous):

anyone know logs? log(d)+log(2d+1)=log(7d)

myininaya (myininaya):

logx+logy=log(xy) so you can use that property here

OpenStudy (anonymous):

so it becomes log(d(2d+1))=log(7d) ? what then

OpenStudy (anonymous):

So you are almost there. Multiply the d through the quantity (2d+1) and according to the property you can set that value equal to 7d and solve for d.

OpenStudy (anonymous):

so log2d^2+d=log7d can just become 2d^2+d=7d?

OpenStudy (anonymous):

that's correct, now solve for d.

OpenStudy (anonymous):

ok thanks :) I never liked logs